OpenStudy (anonymous):

CAN SOMEONE PLEASE HELPP Solve using elimination: -4x = 2 + 2y -7x − 2y = -4 (2, -5) Infinitely many solutions No solution (3, -7)

OpenStudy (anonymous):

Hang on.

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

if your not gonna help dont say hang on

OpenStudy (anonymous):

-4x = 2 + 2y so... -2y = 2 +4x -7x − 2y = -4 so... -7x +2+4x = -4 -3x + 2 = -4 -3x = -6 x = 2 -4(2) = 2 +2y -8 -2 = 2y -10 = 2y y = -5 (2,-5)

OpenStudy (anonymous):

dear suckmath, review this subject, they are easy, you can solve them by yourself

OpenStudy (anonymous):

Thankyou for the help @MattBrotz I get it now(:@tuna theres a reason why my name is suckatmath. Good for you that your good at it though

OpenStudy (anonymous):

no you are not suck at math if you dont want, good luck !

OpenStudy (anonymous):

Not everyone is gonna be good at math so id appreciate that you you have some respect for people who are dumb at it . and im pretty smart at the other subjects btw.

OpenStudy (anonymous):

yes you are right, ok